(pretend exponent ^4)
3. The product of two consecutive integers, n and n + 1, is 42. What is the
positive integer that satisfies the situation?

Answer by stanbon(75887)

(pretend exponent ^4)
3. The product of two consecutive integers, n and n + 1, is 42. What is the
positive integer that satisfies the situation?
----------------------
Equation:
n(n+1) = 42
---
n^2+n-42 = 0
Factor::
(n+7)(n-6) = 0
Positive solution:
n = 6
n+1 = 7
------
Negative solution:
n = -7
n+1 = -6
----------------
Cheers,
Stan H.
-------------------